# The Limits of Price Discrimination Under Privacy Constraints

### v3

@@ Abstract
We consider a producer’s problem of selling a product to a continuum of privacy-conscious consumers, where the producer can implement third-degree price discrimination, offering different prices to different market segments. In the absence of privacy constraints, Bergemann, Brooks, and Morris [2015] characterize the set of all possible consumer-producer utilities, showing that it is a triangle. We consider a privacy mechanism that provides a degree of protection by probabilistically masking each market segment, and we establish that the resultant set of all consumer-producer utilities forms a convex polygon, characterized explicitly as a linear mapping of a certain high-dimensional convex polytope into $\mathbb{R}^{2}$. This characterization enables us to investigate the impact of the privacy mechanism on both producer and consumer utilities. In particular, we establish that the privacy constraint always hurts the producer by reducing both the maximum and minimum utility achievable. From the consumer’s perspective, although the privacy mechanism ensures an increase in the minimum utility compared to the non-private scenario, interestingly, it may reduce the maximum utility. Finally, we demonstrate that increasing the privacy level does not necessarily intensify these effects. For instance, the maximum utility for the producer or the minimum utility for the consumer may exhibit nonmonotonic behavior in response to an increase of the privacy level.
@@ 1 Introduction
1.1 Main results
Our study of the intersection of price discrimination and privacy builds on prior work of Bergemann et al. [2015]. Let us first describe their model and then explain how things change if we consider consumers’ privacy concerns. Consider a monopolist who is selling a product to a continuum of consumers. If the producer knows the values of consumers, then they could engage in first-degree price discrimination to maximize their utility, while the utility of consumers would be reduced to zero. On the other hand, if the producer has no information besides a prior distribution about the value of consumers, then they can only enact uniform pricing. This leads to two pairs of producer utilities and consumer utilities. Bergemann et al. [2015] delineate the set of all possible pairs by using different segmentations. Let us clarify this approach with an example. Imagine a producer aiming to set a product’s price for its customers across the United States. In this context, the aggregated market is defined as the entire US market, meaning the distribution of consumers’ valuations nationwide. Each zip code within the US represents its own market, which is basically the distribution of the consumers’ values in that zip code. Segmentation, in this case, refers to the division of the aggregated US market into these specific markets. Generally speaking, a segmentation involves breaking down the aggregated market into smaller markets, under the condition that these smaller markets must add up to the aggregated market. Bergemann et al. [2015] establish that the set of all possible pairs of consumer and producer utilities is a triangle as depicted in Figure 1. The point $Q$ corresponds to the first-degree price discrimination, and the line $TR$ represents all possible consumer utilities when the producer utility is at its minimum, equating to the utility derived from optimal uniform pricing.
@@ 1.1 Main results
For a given market, knowing the underlying distribution of consumers’ values reveals information about the consumers in that market. Take, for instance, the consumers residing in a certain zip code. Knowing the distribution of the consumers in that zip code induces privacy costs for a multitude of reasons, including potential profiling, harmful targeting, potential manipulation, and unfair business practices. Therefore, we would like to maintain some level of privacy to protect consumers. More precisely, we require that the distribution of consumer values in each market must go through a privacy mechanism that masks the true market and outputs a uniformly sampled market with some probability $\beta$ and outputs the true market with the complementary probability $1-\beta$. Our privacy mechanism is motivated by a market-level randomized response mechanism—where randomized response is a well-known technique in the privacy literature, commonly utilized for the privatization of discrete-valued data [see, e.g., Warner, 1965, Greenberg et al., 1969, Dwork et al., 2014]. Given this framing, we ask the following question:
@@
Notice that for $\beta=0$, our setting reduces to the non-private case, and by increasing $\beta$, we strengthen the privacy guarantees. However, as we now discuss, both our analysis and our results for the setting with privacy constraints are different from those of the non-private case. Specifically, our results allow us to resolve the question of whether the privacy constraint always benefits the consumer. Our main results and insights are as follows.
Main Result 1:
Our point of departure is a setting with two consumer values, a setting in which we are able to fully characterize the space of all consumer and producer utilities. Moreover, this setting allows us to identify nuances that we later explore in a more general setting. In particular, we establish that the set of all consumer and producer utilities is a triangle, similar to the non-private case. However, the triangle takes different shapes depending on the consumer values, the aggregate market distribution, and the privacy parameter. Figure 2 illustrates our characterization for two cases and how it differs from the setting without privacy constraints. In this figure, the set of possible producer and consumer utilities under the privacy mechanism is depicted by the blue triangle $ABC$. The triangle $TQR$, marked by red dashed lines, represents the non-private case for the corresponding set of values and the aggregated market. We identify that the privacy mechanism impacts the limits of price discrimination through a direct effect and an indirect effect. Let us describe these effects and explain how they shape the set of all possible consumer and producer utilities.
@@
Main Result 2:
Our characterization yields the following insights.
@@
Main Result 3:
We establish that our main insights carry over to the general setting with $K>2$ consumer values. Notably, the set of all consumer and producer utilities pairs is no longer a triangle, in sharp contrast to that of the non-private case. Instead, it is a convex polygon (see Figure 3 illustration for one example and how it differs from the non-private case). Similar to the case of $K=2$, we identify that the privacy mechanism introduces both direct and indirect effects on the set of possible consumer and producer utilities. The direct impact implies a combination of scaling and shifting, transforming the triangle $TQR$ into triangle $ABD$ as depicted in Figure 3. The indirect effect of privacy, however, changes the shape of the utility set from a triangle to a convex polygon. More precisely, we illustrate that this set can be represented as a linear mapping of a convex polytope in $\mathbb{R}^{K^{2}}$ into a 2-dimensional space. This representation enables us to extend all the insights above from the case with $K=2$ to the general case with $K\geq 2$. Our analysis also provides the construction of a segmentation to achieve any feasible point for consumer and producer utilities.
@@ 2 Model
where $p_{X}(\cdot)$ and $p_{X|Y}(\cdot|Y)$ are the probability density function and the conditional probability density function of $X$, respectively, and TV(,)normal-$\mathrm{TV}(\cdot,\cdot)$ denotes the total variation distance between two random variables.
@@
It is also worth emphasizing that our privacy mechanism perturbs markets as a whole rather than the values of individual consumers. This means we modify the aggregated data of consumers, not the data of a single consumer. For instance, consider a scenario where market segmentation is based on the zip codes, with each segment representing the market of a single zip code. In this context, our privacy mechanism perturbs the market data for each zip code before it is disclosed to the producer. This approach is particularly fitting in our setting because pricing decisions for a market are based on the distribution of values within that market. Therefore, to ensure a degree of privacy protection, we need to alter the distribution, or essentially, the market itself. Alternatively, if we were to implement a randomized response mechanism at the consumer level, randomly changing each consumer’s value, the continuum of consumers would lead to a deterministic mapping from the actual market to another market. [^4] Even with finite number of consumers, this phenomenon will hold as the number of consumers grows because of the concentration of the realized distribution around its mean. This mapping could be reverse-engineered to the original market, making it ineffective for the pricing problem.
@@
As we discussed above, the producer does not observe the market $x$ and observes a private version $\hat{x}$. Next, we define the optimal pricing of a market given an observation $\hat{x}$ through $\mathcal{M}_{\beta}(\cdot)$. The producer holds a uniform prior over $\mathcal{X}$ and then observes $\hat{x}=\mathcal{M}_{\beta}(x)$. We denote the posterior distribution by $\pi_{\beta}(\cdot|\hat{x})$. It is straightforward to verify that this posterior distribution has a mass of size $1-\beta$ over $\hat{x}$, and the remaining $\beta$ mass is uniformly distributed over $\mathcal{X}$.
@@
$\mathcal{U}_{p}(p|\hat{x}):=\mathbb{E}_{~{}x\sim\pi_{\beta}(x|\hat{x})}\left[\mathcal{U}_{p}(p,x)\right]$ (5)
A price $p$ is called optimal for a market given observation $\hat{x}$ if its corresponding expected revenue is at least as great as that of any other price, i.e.,
@@
Formally, our goal is to characterize the set of all possible pairs of expected utilities for both the producer and consumer across all potential segmentations and optimal pricing rules. In other words, we would like to characterize the following set
@@ 3 Price discrimination limits under privacy: the case of $K=2$
We first focus on the case $K=2$ to gain insights into the limits of price discrimination under privacy constraints. In this scenario, the set of all markets, $\mathcal{X}$, is, in fact, the set of Bernoulli distributions, which can be parameterized by the probability of the high value, $v_{2}$, denoted by $\alpha\in[0,1]$. Throughout this section, we use $x$ and $\alpha$ interchangeably to represent the market. In particular, $\alpha^{*}$ denotes the probability of the high value for the aggregated market $x^{*}$. We also define $\eta$ as the ratio of the lower value to the higher value, i.e.,
$\eta:=\frac{v_{1}}{v_{2}}.$ (9)
We first characterize the optimal price condition on observing $\hat{x}$.
@@
Suppose the producer observes $\hat{x}=(1-\hat{\alpha},\hat{\alpha})$, where $\hat{\alpha}$ denotes the observed probability of the high value $v_{2}$. Let
@@
Then, $v_{1}$ is the optimal price if and only if $\hat{\alpha}\leq t^{*}$, and $v_{2}$ is the optimal price if and only if $\hat{\alpha}\geq t^{*}$.
Note that when $\beta\geq 2\eta$, this lemma implies that the optimal price becomes $v_{2}$, regardless of the observation $\hat{x}$. Similarly, when $\beta\geq 2(1-\eta)$, the optimal price is $v_{1}$. Therefore, if $\beta\geq 2\eta$, then we have
@@
Hence, in what follows, we consider a setting for which the following assumption holds.
@@
In this case, as Lemma 1 establishes, any optimal pricing rule assigns price $v_{1}$ to any market given observation $\hat{\alpha}<t^{*}$ and price $v_{2}$ to any market after observing $\hat{\alpha}>t^{*}$. Next, the subsequent proposition breaks down the impact of the privacy mechanism on the set $\mathcal{S}$, distinguishing between the direct and the indirect effects of privacy.
@@
[…]

### v4

@@ Abstract
We consider a producer’s problem of selling a product to a continuum of privacy-conscious consumers, where the producer can implement third-degree price discrimination, offering different prices to different market segments. Our privacy mechanism provides a degree of protection by probabilistically masking each market segment. We establish that the resultant set of all consumer-producer utilities forms a convex polygon, characterized explicitly as a linear mapping of a certain high-dimensional convex polytope into $\mathbb{R}^{2}$. This characterization enables us to investigate the impact of the privacy mechanism on both producer and consumer utilities. In particular, we establish that the privacy constraint always hurts the producer by reducing both the maximum and minimum utility achievable. From the consumer’s perspective, although the privacy mechanism ensures an increase in the minimum utility compared to the non-private scenario, interestingly, it may reduce the maximum utility. Finally, we demonstrate that increasing the privacy level does not necessarily intensify these effects. For instance, the maximum utility for the producer or the minimum utility for the consumer may exhibit nonmonotonic behavior in response to an increase of the privacy level.
@@ 1 Introduction
Our study of the intersection of price discrimination and privacy builds on prior work of Bergemann et al. [2015]. Let us first describe their model and then explain how things change if we consider consumers’ privacy concerns. Consider a monopolist who is selling a product to a continuum of consumers. If the producer knows the values of consumers, then they could engage in first-degree price discrimination to maximize their utility, while the utility of consumers would be reduced to zero. On the other hand, if the producer has no information besides a prior distribution about the value of consumers, then they can only enact uniform pricing. This leads to two pairs of producer utilities and consumer utilities. Bergemann et al. [2015] delineate the set of all possible pairs by using different segmentations. Let us clarify this approach with an example. Imagine a producer aiming to set a product’s price for its customers across the United States. In this context, the aggregated market is defined as the entire US market, meaning the distribution of consumers’ valuations nationwide. Each zip code within the US represents its own market, which is basically the distribution of the consumers’ values in that zip code. Segmentation, in this case, refers to the division of the aggregated US market into these specific markets. Generally speaking, a segmentation involves breaking down the aggregated market into smaller markets, under the condition that these smaller markets must add up to the aggregated market. Bergemann et al. [2015] establish that the set of all possible pairs of consumer and producer utilities is a triangle as depicted in Figure 1. The point $Q$ corresponds to the first-degree price discrimination, and the line $TR$ represents all possible consumer utilities when the producer utility is at its minimum, equating to the utility derived from the uniform pricing.
@@
For a given market, knowing the underlying distribution of consumers’ values reveals information about the consumers in that market. Take, for instance, the consumers residing in a certain zip code. Knowing the distribution of the consumers in that zip code induces privacy costs for a multitude of reasons, including potential profiling, harmful targeting, potential manipulation, and unfair business practices. In this paper, we aim to explore how limiting information disclosure about the market affects the potential combinations of consumer and seller utilities. We conceptualize privacy leakage as the extent to which revealing a message about the value distribution within a market segment reveals the actual distribution of values of consumers in that segment. This measure quantifies the change in our understanding of a segment’s distribution following the receipt of a specific message. More precisely, we require that the distribution of consumer values in each market must go through a privacy mechanism that masks the true market and outputs a uniformly sampled market with some probability $\beta$ and outputs the true market with the complementary probability $1-\beta$. Our privacy mechanism is motivated by a market-level randomized response mechanism—where randomized response is a well-known technique in the privacy literature, commonly utilized for the privatization of discrete-valued data [see, e.g., Warner, 1965, Greenberg et al., 1969]. Additionally, as we will highlight in the next section when presenting our model, our privacy mechanism is methodology used in practice in technology companies and government agencies when aggregate statistics such as histograms are published in a manner that provides privacy guarantees Desfontaines [2021]. These methods ensure that a function reported over users’ data does not significantly depend on any individual user’s data, thereby safeguarding personal privacy Dwork et al. [2014].
In this paper, we operate within a Bayesian setting where we assume that the producer holds a prior for every market and updates it after observing the output of the privacy mechanism. To avoid any unaccounted privacy loss, it is important to choose the least informative prior. Intuitively, under the principle of indifference, a reasonable choice of prior belief for the producer would be the uniform distribution over the entire space of market distributions. Indeed, the question of selecting an uninformative prior has been the subject of extensive literature in Bayesian analysis, where reference priors are a leading framework Bernardo [1979]. As we will demonstrate, the uniform distribution is the reference prior in our setting. Given this Bayesian framework, we ask the following question:
@@
Notice that by increasing $\beta$, we strengthen the privacy guarantees. The case of $\beta=1$ corresponds to a fully private setting in which the producer has no information about each segment or the aggregated market. It is worth emphasizing that in our setting, unlike in Bergemann et al. [2015], the producer does not know the aggregated market. This assumption is crucial when studying privacy loss, as having additional knowledge about the aggregated market would constitute unaccounted information leakage. Consider, for example, a setup with two possible values where each consumer’s value is either low or high. If the aggregated market indicates that everyone’s value is high with a probability of one, then the producer would know every segment’s market, even if $\beta=1$ (which is supposed to deliver a full privacy guarantee). Lastly, the case of $\beta=0$ reduces our setting to the non-private case, where the producer can see all segments without any perturbation, and hence can deduce the aggregated market as well.
Both our analysis and our results are different in significant ways from those of the non-private case. In particular, since the aggregated market is no longer known to the producer, there is no guarantee that the producer’s surplus is at least as high as profits under the uniform monopoly price (i.e., at or above the line $TR$ in Figure 1). Moreover, our results allow us to resolve the question of whether the privacy constraint always benefits the consumer. In more detail, our main results and insights are as follows.
First, in the setting of two consumer values, we are able to fully characterize the space of all consumer and producer utilities. Moreover, we are able to identify nuances in this setting that we later explore in a more general setting. In particular, we establish that the set of all consumer and producer utilities is a triangle, similar to the non-private case, but the triangle takes different shapes depending on the consumer values, the aggregate market distribution, and the privacy parameter. Figure 2 illustrates our characterization for two cases that display qualitative differences compared to the setting without privacy constraints. In this figure, the set of possible producer and consumer utilities under the privacy mechanism is depicted by the blue triangle $ABC$. The triangle $TQR$, marked by red dashed lines, represents the non-private case for the corresponding set of values and the aggregated market. We see that the privacy mechanism impacts the limits of price discrimination through a direct effect and an indirect effect. Let us describe these effects and explain how they shape the set of all possible consumer and producer utilities.
@@
Second, our characterization yields the following insights.
@@
Third, we establish that our main insights carry over to the general setting with $K>2$ consumer values. Notably, the set of all consumer and producer utilities pairs is no longer a triangle, in sharp contrast to that of the non-private case. Instead, it is a convex polygon (see Figure 3 illustration for one example and how it differs from the non-private case). Similar to the case of $K=2$, we identify that the privacy mechanism introduces both direct and indirect effects on the set of possible consumer and producer utilities. The direct impact implies a combination of scaling and shifting, transforming the triangle $TQR$ into triangle $ABD$ as depicted in Figure 3. The indirect effect of privacy, however, changes the shape of the utility set from a triangle to a convex polygon. More precisely, we illustrate that this set can be represented as a linear mapping of a convex polytope in $\mathbb{R}^{K^{2}}$ into a 2-dimensional space. This representation enables us to extend all the insights above from the case with $K=2$ to the general case with $K\geq 2$. Our analysis also provides the construction of a segmentation to achieve any feasible point for consumer and producer utilities.
@@ 2 Model
where $p_{X}(\cdot)$ and $p_{X|Y}(\cdot|Y)$ are the probability density function and the conditional probability density function of $X$, respectively, and $\mathrm{TV}(\cdot,\cdot)$ denotes the total variation distance between two random variables.
@@
Our privacy mechanism privatizes the distribution of values, i.e., the market, and is therefore closely related to the literature on differentially private histograms, with applications ranging from limiting inferences about the prevalence of a disease in a group of people to protecting password frequency lists (see for instance Xu et al. [2013], Suresh [2019]), and with fielded applications in companies and agencies such as Apple, Google, and the United States Census Bureau Apple , Abowd [2018]. Briefly, a randomized algorithm, as a function of users’ data, is considered differentially private if the distribution of the algorithm’s output does not change significantly by altering the data of one user. The motivation behind this definition is that if the output is not overly sensitive to a user’s data, it implies that we cannot learn much about a user by observing the output [Dwork et al., 2006a, b]. In Section A.2, we further discuss differential privacy and its connection to the privacy mechanism in (3).
@@
Producer’s belief update:
As we discussed in the introduction, the producer does not know the aggregated market or any of the segment markets. Instead, they hold a prior regarding each segment’s market and then update it using Bayes’ theorem after observing the market through the privacy mechanism. We denote the prior by $\pi(\cdot)$ and the posterior after observing $\hat{x}=\mathcal{M}_{\beta}(x)$ by $\pi_{\beta}(\cdot|\hat{x})$. Note that both distributions are defined over the space of all markets $\mathcal{X}$.
Here, we need to make a modeling assumption about the choice of prior. In particular, we want the prior to carry no information about the underlying market, as otherwise, it would mean a privacy leakage that is not accounted for. Consider, for instance, a setting with $K=2$ values where all consumers have the higher value ($v_{2}$), i.e., $x^{*}(v_{2})=1$. If the prior assigns a high probability to the markets where the probability of the high value is close to one, i.e., $x$ such that $x(v_{2})\approx 1$, then even with $\beta=1$, the posterior would still place a high probability on markets with $x(v_{2})$ close to one. This implies that the seller almost knows that most of the market comprises high-value consumers. In other words, even $\beta=1$ fails to provide strong privacy protection. In the remainder of this paper we make the following assumption concerning the prior distribution $\pi(\cdot)$.
Assumption 1.
The prior $\pi(\cdot)$ is equal to $\text{Unif}(\mathcal{X})$ where $\text{Unif}(\mathcal{X})$ denotes the uniform distribution over the space of all markets $\mathcal{X}$. In other words, $\pi(x)=\tfrac{1}{\text{Vol}(\mathcal{X})}$ for any market $x\in\mathcal{X}$.
This assumption ensures that we select a least informative prior, which can be formalized variationally as the prior that maximizes the expected Kullback-Leibler (KL) divergence of posterior from the prior [Bernardo, 1979].
Definition 2.
The reference prior is the prior that maximizes the KL divergence in expectation:
[…]
where $\mu_{\beta}(\cdot)$ denotes the distribution of the observed market $\hat{x}$
The reference prior in our setting is the uniform distribution over the space of markets $\mathcal{X}$.
Proposition 2.
$\text{Unif}(\mathcal{X})$ is the solution to (5).
In fact, if we want to choose a prior that conveys no information about the underlying market, the prior distribution is an intuitive choice, as all different markets are equally likely under such a prior. Therefore, the producer starts with $\text{Unif}(\mathcal{X})$ as their prior belief, and after observing $\hat{x}=\mathcal{M}_{\beta}(x)$, they form the posterior. By applying Bayes’ theorem, it can be demonstrated that this posterior distribution has a mass of size $1-\beta$ over $\hat{x}$, and the remaining $\beta$ mass is uniformly distributed over $\mathcal{X}$.
@@
$\mathcal{U}_{p}(p|\hat{x}):=\mathbb{E}_{~{}x\sim\pi_{\beta}(x|\hat{x})}\left[\mathcal{U}_{p}(p,x)\right].$ (6)
A price $p$ is called optimal for a market given observation $\hat{x}$ if its corresponding expected revenue is at least as great as that of any other price:
@@
It is worth highlighting that the above utilities are computed with respect to the true market $x$ (and not as an expectation with respect to the prior $\text{Unif}(\mathcal{X})$). In other words, there is a fixed market $x$, and the producer has a uniform prior $\text{Unif}(\mathcal{X})$ on it. This prior is updated to the posterior $\pi_{\beta}(x|\hat{x})$ after observing $\hat{x}\sim\mathcal{M}_{\beta}(x)$. Based on this posterior, the producer selects the optimal price $p\sim\phi(\hat{x})$. Now, the functions $U_{p}^{\phi}(x)$ and $U_{c}^{\phi}(x)$ calculate the utilities for the producer and consumer of the original market $x$ when priced at $p$.
Formally, our goal is to characterize the set of all possible pairs of expected utilities for both the producer and consumer across all potential segmentations and optimal pricing rules. In other words, we would like to characterize the following set:
@@ 3 Price discrimination limits under privacy: the case of $K=2$
We begin by examining the simpler scenario where the consumer’s valuation is limited to two values, i.e., $K=2$, to gain insights into the limits of price discrimination under privacy constraints. In this scenario, the set of all markets, $\mathcal{X}$, is, in fact, the set of Bernoulli distributions, which can be parameterized by the probability of the high value, $v_{2}$, denoted by $\alpha\in[0,1]$. Throughout this section, we use $x$ and $\alpha$ interchangeably to represent the market. In particular, $\alpha^{*}$ denotes the probability of the high value for the aggregated market $x^{*}$. We also define $\eta$ as the ratio of the lower value to the higher value, i.e.,
$\eta:=\frac{v_{1}}{v_{2}}\leq 1.$ (10)
We first characterize the optimal price conditioning on the observed market $\hat{x}$.
@@
Suppose the producer observes $\hat{x}=(1-\hat{\alpha},\hat{\alpha})$, where $\hat{\alpha}$ denotes the observed probability of the high value $v_{2}$. Define the threshold $t^{*}$ as:
@@
Then, the producer should set the price to $v_{1}$ (i.e., it is the optimal price) if and only if $\hat{\alpha}\leq t^{*}$, and $v_{2}$ is the optimal price otherwise.
Note that when $2\eta\leq\beta$, this lemma implies that the optimal price becomes $v_{2}$, regardless of the observation $\hat{x}$. Similarly, when $\beta\geq 2(1-\eta)$, the optimal price is $v_{1}$. Note that when the producer sets the price to the high value $v_{2}$, only consumers with the high value will purchase the product, resulting in their expected utility being zero due to the absence of consumer surplus. Conversely, when the producer opts for the low value $v_{1}$ as the price, all consumers—regardless of their valuation—will purchase the product. This pricing strategy allows consumers with the high value to achieve a positive expected surplus, as they gain the difference between their valuation and the lower price. Therefore, if $\beta\geq 2\eta$, we have
@@
We next focus on the more interesting scenario in which the following assumption holds.
@@
As established in Lemma 1, in this case, the optimal pricing rule sets the price to $v_{1}$ for observations $\hat{\alpha}<t^{*}$ and $v_{2}$ for $\hat{\alpha}>t^{*}$. The following proposition analyzes the effects of privacy mechanism on the set $\mathcal{S}$, distinguishing between the direct and the indirect effects of privacy.
@@
$\qquad\alpha_{1}\in[0,\min\{t^{*},\alpha^{*}\}],\alpha_{2}\in[\max\{t^{*},\alpha^{*}\},1],\gamma_{1}\text{ and }\gamma_{2}\in[0,1],\gamma_{1}+\gamma_{2}=1,\gamma_{1}\alpha_{1}+\gamma_{2}\alpha_{2}=\alpha^{*}\biggr{\}}.$